# Bifurcation of time crystals in driven and dissipative Rydberg atomic gas

### 22 Feb 2024 v2

@@
A time crystal is an exotic phase of matter where time-translational symmetry is broken; this phase differs from the spatial symmetry breaking induced in crystals in space. Lots of experiments report the transition from a thermal equilibrium phase to time crystal phase. However, there is no experimental method to probe the bifurcation effect of distinct time crystals in quantum many-body systems. Here, in a driven and dissipative many-body Rydberg atom system, we observe multiple continuous dissipative time crystals and emergence of more complex temporal symmetries beyond the single time crystal phase. Bifurcation of time crystals in strongly interacting Rydberg atoms is observed; the process manifests as a transition from a time crystal state of long temporal order to one of short temporal order, or vice versa. By manipulating the driving field parameters, we observe the time crystal’s bistability and a hysteresis loop. These investigations indicate new possibilities for control and manipulation of the temporal symmetries of non-equilibrium systems.
@@
The term time crystal refers to a unique state of matter in which a system exhibits spontaneous breaking of its time translation symmetry; this state was initially proposed by Frank Wilczek [1]. In other words, the time crystal is a phase of matter in which the system’s behavior repeats periodically in both space and time. A time crystal exhibits long-range order in time, whereas conventional crystals exhibit long-range order in space. The state includes discrete and continuous time crystal phases [2, 3, 4, 5] that exhibit discrete and continuous time-translation symmetry breaking, respectively. Since the initial proposal, time crystals were first reported in experiments using trapped ions [6] and the nitrogen-vacancy centers in diamonds [7]. There has been considerable progress on both the theoretical [8, 9, 10, 11, 12] and experimental fronts in efforts to study and understand time crystals [13, 14, 15, 16, 17, 18, 19]. Researchers continue to strive to improve the stability [20] and the coherence [21, 22] of time crystals, and to probe the equilibrium of quantum matter [23]. Interestingly, the system comes out of thermal equilibrium dissipation under external driving conditions, and researchers have observed stabilized dissipative time crystals [24], prethermal discrete time crystals [25, 26], and higher-order time crystals [27, 28].
@@
The large dipole moment of the Rydberg atom provides a good platform for study of the dynamics of strongly correlated systems. For example, researchers have used Rydberg atoms to probe quantum phase transitions [29, 30], quantum scars [31, 32], non-equilibrium phase transitions, and self-organized criticality [33, 34, 35, 36, 37, 38]. In traditional equilibrium systems, the second law of thermodynamics dictates that entropy will always increase with time, and this leads to thermal equilibrium and the absence of persistent oscillatory behavior. In the driven and dissipative Rydberg system, however, the external driving force injects energy into the system, while the long-range interactions between the system’s constituents aid in redistributing this energy and thus help to maintain a persistent oscillatory evolution of the Rydberg atom population [39, 40, 41, 42]. The driven and dissipative Rydberg atom system is inherently an out-of-equilibrium complex system, which means that it does not settle into a steady state or thermal equilibrium, and the exotic emergence and transition dynamics of time crystals are worthy of study.
@@ Acknowledgements
We acknowledge funding from the National Key R and D Program of China (Grant No. 2022YFA1404002), the National Natural Science Foundation of China (Grant Nos. U20A20218, 61525504, and 61435011), the Anhui Initiative in Quantum Information Technologies (Grant No. AHY020200), and the Major Science and Technology Projects in Anhui Province (Grant No. 202203a13010001).

### 27 Feb 2024 v3

@@
A time crystal is an exotic phase of matter where time-translational symmetry is broken; this phase differs from the spatial symmetry breaking induced in crystals in space. Lots of experiments report the transition from a thermal equilibrium phase to time crystal phase. However, there is no experimental method to probe the bifurcation effect of distinct continuous time crystals in quantum many-body systems. Here, in a driven and dissipative many-body Rydberg atom system, we observe multiple continuous dissipative time crystals and emergence of more complex temporal symmetries beyond the single time crystal phase. Bifurcation of time crystals in strongly interacting Rydberg atoms is observed; the process manifests as a transition from a time crystal state of long temporal order to one of short temporal order, or vice versa. By manipulating the driving field parameters, we observe the time crystal’s bistability and a hysteresis loop. These investigations indicate new possibilities for control and manipulation of the temporal symmetries of non-equilibrium systems.
@@
The term time crystal refers to a unique state of matter in which a system exhibits spontaneous breaking of its time translation symmetry; this state was initially proposed by Frank Wilczek [1]. In other words, the time crystal is a phase of matter in which the system’s behavior repeats periodically in both space and time. A time crystal exhibits long-range order in time, whereas conventional crystals exhibit long-range order in space. The state includes discrete and continuous time crystal phases [2, 3, 4, 5] that exhibit discrete and continuous time-translation symmetry breaking, respectively. Since the initial proposal, time crystals were first reported in experiments using trapped ions [6] and the nitrogen-vacancy centers in diamonds [7]. There has been considerable progress on both the theoretical [8, 9, 10, 11, 12, 13, 14] and experimental fronts in efforts to study and understand time crystals [15, 16, 17, 18, 19, 20, 21]. Researchers continue to strive to improve the stability [22] and the coherence [23, 24] of time crystals, and to probe the equilibrium of quantum matter [25]. Interestingly, the system comes out of thermal equilibrium dissipation under external driving conditions, and researchers have observed stabilized dissipative time crystals [26], prethermal discrete time crystals [27, 28], and higher-order time crystals [29, 30].
@@
The large dipole moment of the Rydberg atom provides a good platform for study of the dynamics of strongly correlated systems. For example, researchers have used Rydberg atoms to probe quantum phase transitions [31, 32], quantum scars [33, 34], non-equilibrium phase transitions, and self-organized criticality [35, 36, 37, 38, 39, 40]. In traditional equilibrium systems, the second law of thermodynamics dictates that entropy will always increase with time, and this leads to thermal equilibrium and the absence of persistent oscillatory behavior. In the driven and dissipative Rydberg system, however, the external driving force injects energy into the system, while the long-range interactions between the system’s constituents aid in redistributing this energy and thus help to maintain a persistent oscillatory evolution of the Rydberg atom population [41, 42], and generating dissipative continuous time crystal [43]. The driven and dissipative Rydberg atom system is inherently an out-of-equilibrium complex system, which means that it does not settle into a steady state or thermal equilibrium, and the exotic emergence and transition dynamics of distinct continuous time crystals are worthy of study.
@@ Discussion
Note. While finishing this manuscript we became aware of a related work of reporting a phase transition from a continuous to a discrete time crystal [Ref. [47]], and a related work (Ref. [48]) observing time crystal comb.
@@ Acknowledgements
D.-S.D. thanks for the previous discussions with Prof. Igor Lesanovsky and Prof. C. Stuart Adams on time crystals. We acknowledge funding from the National Key R and D Program of China (Grant No. 2022YFA1404002), the National Natural Science Foundation of China (Grant Nos. U20A20218, 61525504, and 61435011), the Anhui Initiative in Quantum Information Technologies (Grant No. AHY020200), and the Major Science and Technology Projects in Anhui Province (Grant No. 202203a13010001).